OpenStudy (anonymous):

The cost to produce a custom designed t-shirt is five dollars per shirt plus a 40 dollar design fee. Which of the following could be data points that would fit this situation? A (1,40)AND(2,45) B (45,1)AND(50,2) C (1,45)AND(2,50) D (1,40)AND(2,50)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

c i tshirt =5+40 =45 (1,45) 2 tshirts = 5x2 + 40 = 10 + 40 = 50 (2,50)